40..0ml of 0.400M NaOH solution is diluted to a final voulem o 200.0 mL, calculate the new concentration
1
Expert's answer
2012-10-30T10:04:43-0400

V1= 40 mL
C1=0.4 M
V2= 200 mL
C2 - ?
V1*C1 = V2*C2  C2 = (V1*C1)/V2

C2 = (40mL*0.4M)/200mL = 0.08 M

C2 = 0.08 M